OpenStudy (anonymous):

What is the period of the function? y=2 cos pi x

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

period of \(\cos(bx)\) is \(\frac{2\pi}{b}\) in your case \(b=\pi\) so the period is \(\frac{2\pi}{\pi}=2\)

OpenStudy (anonymous):

A man stands on his balcony, 130 feet \above the ground. He looks at the ground, with his sight line forming an \angle of 70 degrees with the building, and sees a bus stop. The function d=130 \sec \theta \models the distance from the man \to any object given his \angle of sight \theta. How far is the bus stop from the man? Round \to your nearest foot

OpenStudy (anonymous):

@satellite73 @iambatman

OpenStudy (anonymous):

did you ever get the answer